OpenStudy (anonymous):

5. Write the equation for the vertical line that contains point E(–7, 7). (1 point) y = 7 x = –7 y = –7 x = 7

OpenStudy (anonymous):

Please, at least try to do it yourself.

OpenStudy (anonymous):

It's very simple once you try.

OpenStudy (anonymous):

Very simple to you. Not me.

OpenStudy (anonymous):

Okay, sorry. But can you ask your teacher to explain it better at least? I'm sure they would be glad to help if you asked.

OpenStudy (anonymous):

I have. They are busy with the other students.

OpenStudy (anonymous):

So "c"?

OpenStudy (anonymous):

Hm. I see you don't have a good teacher? Okay. I'll use another example for you and see if you get it.

OpenStudy (anonymous):

Write the equation for the vertical line that contains point E(10, –3).? Any straight line can be written with the equation y=mx+b, where m is the slope of the line (rise/run) and b is where the line intersects the y axis. In this particular case, your line crosses the x-axis at 10 and crosses the y-axis at -3. So using the process above, we find the equation to be y=(3/10)x-3

OpenStudy (anonymous):

I'll just guess... I do not understand this...

OpenStudy (anonymous):

Hm. That will not help you. If you look at the problem, just subsitute your own numbers into the problem and solve it from there.

OpenStudy (anonymous):

Write the equation for the vertical line that contains point E(-7,7).? Any straight line can be written with the equation y=mx+b, where m is the slope of the line (rise/run) and b is where the line intersects the y axis. In this particular case, your line crosses the x-axis at -7 and crosses the y-axis at 7. So using the process above, we find the equation to be y=(-7/7)x-7

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

Do you understand better now?

OpenStudy (anonymous):

Yes, Thank you.

OpenStudy (anonymous):

No problem. Sorry for being a bit rude before.

OpenStudy (anonymous):

It's cool.

OpenStudy (anonymous):

Okay, great!